# Decipher-MR: A Vision-Language Foundation Model for 3D MRI Representations

### v1

@@ Abstract
Magnetic Resonance Imaging (MRI) is a critical medical imaging modality in clinical diagnosis and research, yet its complexity and heterogeneity pose challenges for automated analysis, particularly in scalable and generalizable machine learning applications. While foundation models have revolutionized natural language and vision tasks, their application to MRI remains limited due to data scarcity and narrow anatomical focus. In this work, we present Decipher-MR, a 3D MRI-specific vision-language foundation model trained on a large-scale dataset comprising 200,000 MRI series from over 22,000 studies spanning diverse anatomical regions, sequences, and pathologies. Decipher-MR integrates self-supervised vision learning with report-guided text supervision to build robust, generalizable representations, enabling effective adaptation across broad applications. To enable robust and diverse clinical tasks with minimal computational overhead, Decipher-MR supports a modular design that enables tuning of lightweight, task-specific decoders attached to a frozen pretrained encoder. Following this setting, we evaluate Decipher-MR across diverse benchmarks including disease classification, demographic prediction, anatomical localization, and cross-modal retrieval, demonstrating consistent performance gains over existing foundation models and task-specific approaches. Our results establish Decipher-MR as a scalable and versatile foundation for MRI-based AI, facilitating efficient development across clinical and research domains.
@@ 2.1 Decipher-MR Pretraining
Decipher-MR was pretrained on a large and diverse MRI dataset comprising 22,594 studies and 203,233 series (Fig. 1a), with corresponding radiology reports available for 20,658 studies. The dataset spans a broad age range (0 to 90 years), covers a wide spectrum of body regions and MRI protocols, and features scans acquired from multiple vendors, including GE, Siemens, Philips, Toshiba, etc. This diversity ensures robust representation across various imaging systems and clinical environments.
@@ 2.2 Decipher-MR Learns Robust Visual Representation for Classification Tasks
A broad range of clinical and radiological tasks involving MRI can be formulated as classification problems, including disease diagnosis, demographic prediction, body region identification, image sequence classification, and contrast detection. We benchmarked our method against state-of-the-art foundation models using a simple probing setup, where the pretrained encoder was frozen and only a lightweight three-layer MLP attached on top was fine-tuned across these varied tasks. Our model consistently outperformed DINOv2[4] and BiomedCLIP[14] by a large margin, and demonstrated clear advantages over MedImageInsight[12], achieving average gains of 2.9% in disease classification, 3.0% in demographic prediction, and 0.2% in imaging-related attribute detection (Fig. 2a). Performance differences between models became more pronounced under limited data settings (Supplementary Table 3-5), especially in tasks such as imaging attribute detection where performance otherwise tends to saturate (Fig. 2c). Further details on the classification tasks within each category, along with the datasets used, are provided in Methods Section 4.8 and Supplementary Table 2.
@@
These findings highlight the strength of our model in capturing transferable and clinically meaningful imaging features, with the potential to improve diagnostic accuracy, enhance clinical workflows, and reduce errors. Importantly, most of evaluations were performed without task-specific preprocessing (e.g., skull stripping, registration, or organ cropping) or fine-tuning, ensuring fair and direct model comparisons. Further performance improvements are possible through specialized model architectures and task-optimized data preprocessing. For instance, simply resizing MR images in-plane to match MedImageInsight’s input resolution substantially improved performance in tasks such as liver and prostate lesion classification, brain age regression, and attribute detection (Fig. 2a). Additional cropping around the region of interest can further enhance performance for localized tasks—for instance, prostate lesion classification improved by 7% when cropping around the prostate (ProstateCrop vs. Prostate in Fig.2a and Supplementary Table 3), further amplifying Decipher-MR’s advantage over other models. However, such cropping can impair performance on tasks that rely on broader anatomical context, such as age prediction from prostate MRI (Supplementary Table 3), which benefits from information beyond the prostate region.
@@ 2.3 Decipher-MR Benefits from Text and Diversity in Pretraining Data
Significant performance gains were observed with second-stage image–text contrastive pretraining, which consistently outperformed models trained solely on imaging data (Fig. 2b). These results highlight the benefit of weak supervision from radiology reports and demonstrate the effectiveness of our two-stage pretraining strategy. Notable improvements were found in cardiac disease classification (+5.0%), prostate lesion classification (+2.4%), and several low-data scenarios, including brain age regression (+3.2%), body region detection (+4.2%), and contrast detection (+1.6%), with the latter two particularly pronounced under low-data scenarios (Fig. 2c).
@@ 2.4 Decipher-MR Supports Cross-modality Retrieval
On in-domain datasets, querying with either full reports or just the conclusion sections (Supplementary Fig. 1) retrieved the correct paired scan within the top 10 results in approximately 26% of cases, significantly outperforming MedImageInsight (5.1%) (Supplementary Table 6).
@@ 2.5 Decipher-MR Enables Robust Medical Image Segmentation with Fast Convergence
Across all datasets, Decipher-MR paired with a ResNet-based decoder (Method 4.10) consistently outperformed MedSAM3D[24] under the same frozen-encoder setting. Remarkably, even though only the decoder was trained, the performance was comparable to nnUNet[25], a widely used, fully tuned, state-of-the-art end-to-end segmentation framework in medical imaging (Fig. 4a). Qualitative examples in Fig. 4c show accurate segmentation of small abdominal organs and cardiac structures at both end-diastole (ED) and end-systole (ES), closely aligned with ground truth annotations. These results underscore the strong representational power of Decipher-MR across diverse anatomical regions.
@@ 2.6 Decipher-MR Supports Anomaly Localization and Visual Grounding
Beyond classification and organ segmentation, accurate pathology or anomaly localization is critical for the clinical deployment of AI models, especially when integrating human input. To investigate this, we evaluated Decipher-MR on anomaly localization tasks involving head and neck tumors and surgically removed organs (e.g., prostatectomy, hepatectomy, and pneumonectomy) (Fig. 4e), under two settings: standalone localization and visual grounding guided by human-provided text input describing anatomical location and pathology types. For each setting, we adopted state-of-the-art task-specific frameworks: DETR3D[26] for localization and MedRPG[27] for visual grounding, replacing and initializing their visual encoders with Decipher-MR (Method 4.11).
@@ 3 Discussion
In this study, we introduced Decipher-MR, an MRI-specific 3D vision-language foundation model pretrained on a large and diverse collection of MR datasets. We evaluated its capabilities across a comprehensive range of downstream tasks—including classification, cross-modal retrieval, organ segmentation, and anomaly localization—spanning multiple stages of the clinical workflow. Experimental results on both propriety and public datasets demonstrated that Decipher-MR is a powerful feature extractor for MR images. It consistently outperformed other foundation models on simple probing tasks and achieved performance on par with or exceeding state-of-the-art, task-specific models trained end-to-end. One particularly interesting finding is that Decipher-MR maintains strong performance even though its pretraining data primarily consists of individuals under the age of 70 and rarely include cases of Alzheimer’s disease [28]. Despite this, it performs competitively on related tasks, suggesting a degree of generalization beyond the original pretraining distribution.
@@
Despite these advantages, several limitations remain. While Decipher-MR showed strong performance in cross-modal retrieval, particularly for anatomical region retrieval, its effectiveness in pathology-focused retrieval was more limited. This likely stems from insufficient diversity in textual report data during pretraining. Additionally, we did not observe significant improvements in imaging attribute classification (e.g., sequence or contrast type) using additional text supervision. Future work could incorporate richer DICOM metadata during pretraining to improve model capability in these areas and enable more radiology-relevant search functionalities. Finally, while our frozen encoder setup has shown robust performance, encoder fine-tuning may be necessary to unlock further gains for certain applications. We plan to explore various tuning strategies to optimize the trade-off between efficiency and task-specific performance.
@@ 4.2 Evaluation Dataset
Validation set of our pretraining datasets: Contains around 2,000 studies and 20,000 MRI series, with radiology reports for most studies. MRI sequence details, body regions, and contrast usage were extracted from DICOM headers.
@@
ADNI: A dataset for Alzheimer’s disease research, containing diagnoses for cognitively normal (CN), mild cognitive impairment (MCI), and Alzheimer’s disease (AD) individuals, along with patient demographics. We used 1,700 baseline T1-weighted MRIs from ADNI1 and ADNI2 for classification evaluations.[36]
PICAI: A prostate cancer dataset comprising 1,500 bi-parametric MRI (bpMRI) exams per patient, including T2-weighted (T2W), high b-value diffusion-weighted imaging (DWI), and apparent diffusion coefficient (ADC) maps, along with annotated prostate and tumor masks and diagnostic labels (benign vs. malignant) [37].
ACDC: A cardiac MRI dataset with 150 exams across five subgroups, including four pathological conditions (myocardial infarction, dilated cardiomyopathy, hypertrophic cardiomyopathy, abnormal right ventricle) and one healthy group, along with segmentation masks for three heart regions [23].
LLD-MMRI: A multi-phase liver MRI dataset for lesion diagnosis, covering seven lesion types: three malignant (intrahepatic cholangiocarcinoma, liver metastases, hepatocellular carcinoma) and four benign (hepatic hemangioma, hepatic abscess, hepatic cysts, focal nodular hyperplasia) [38].
MRART: A dataset of T1-weighted structural MRI head scans capturing three levels of motion artifacts [39].
@@ 4.3 Model Architecture
For Decipher-MR, we employed a 3D Vision Transformer (ViT) [40] as the image encoder, using non-overlapping 8×8×8 volumetric patches. Each patch was flattened and passed through a linear projection to generate patch embeddings. The rest of the architecture follows the standard ViT-Base configuration, consisting of 86 million parameters and a 768-dimensional embedding size. Trilinear position embedding interpolation enables the model to handle inputs of variable sizes during inference and downstream applications.
@@ 4.4 Model Pretraining
4.4.1 First stage (vision): Student-Teacher Self-Supervised Learning for Vision Encoder
@@
4.4.2 First stage (text): Masked Language Pretraining for Text Encoder
@@
4.4.3 Second stage: Image-Report Contrastive Learning
In the second stage of pretraining, we leveraged the pretrained vision and text encoders from the first stage and further refined them using an Image-Report Contrastive Learning approach [3] to align image and text embeddings in a shared multi-modal embedding space. During training, MRI-report pairs are processed to generate their respective embeddings. The model encourages embeddings from matching image-report pairs to be close in the shared space while pushing apart embeddings from non-matching pairs. Details regarding the construction of image and text inputs, along with the data sampling strategy designed to enhance contrastive learning, are provided in Section 4.4.5.
4.4.4 Data Augmentation
To enable a generalist foundation model capable of handling inconsistencies in MRI resolution, size, and acquisition protocols across scanners and body regions, we tailored our vision encoder training to be robust to such inconsistencies. Beyond standard augmentations like rotation and flipping, we customized the cropping strategy used in DINO’s teacher network. DINOv2 includes two global crops, and we adapted them for MRI by defining one crop that covers the entire scan resized to the input dimensions, and another partial crop extracted at a randomly selected resolution to introduce scale diversity. From these two global crops, we generated local crops for the student network, enabling the model to process the same scan at multiple resolutions and learn consistent image-level representations. In the second stage of multi-modal contrastive training, we preserved more complete contextual information while keeping robustness to resolution variation by randomly cropping within 75%–100% of the image size and resizing to a relatively larger input size to retain fine details (Method 4.5).
4.4.5 Organ-Based Study-level Data Sampling
@@ 4.5 Pretraining Data Processing
MRI Processing:
We first applied foreground cropping to all MRI images using Otsu’s method to remove the black background. Each image was then rescaled to the range of -1 to 1 based on the 1st and 99th percentiles of pixel intensities. For the first-stage pretraining, we set the global crop size to 256×256×24 and the local crop size to 128×128×16, selected based on the median dimensions of our pretraining dataset. As described in Section 4.4.4, the first global crop was resized to match the target size while preserving the x-y aspect ratio, with black padding added as needed The second global crop was a random crop with variable resolution. In the second-stage contrastive training, the larger crops mentioned in Section 4.4.4 were resized to 384×384×48 to provide a more comprehensive view within each crop.
Report Processing: We first used Llama 3.3 [43] to process the noisy MRI reports by removing study details unrelated to imaging and extracting organ-specific descriptions and conclusions into separate sections, including any identified abnormalities. We then mapped each organ section to a list of standard SNOMED anatomical entities [44] using AWS Comprehend Medical [45] for the sampling process described in Section 4.4.4. Additional details, including the prompt used and examples of reports before and after processing, are provided in Supplementary Fig. 1.

### v2

@@ Abstract
Magnetic Resonance Imaging is a critical imaging modality in clinical diagnosis and research, yet its complexity and heterogeneity hinder scalable, generalizable machine learning. Although foundation models have revolutionized language and vision tasks, their application to MRI remains constrained by data scarcity and narrow anatomical focus. We present Decipher-MR, a 3D MRI-specific vision-language foundation model trained on 200,000 MRI series from over 22,000 studies spanning diverse anatomical regions, sequences, and pathologies. Decipher-MR integrates self-supervised vision learning with report-guided text supervision to build robust representations for broad applications. To enable efficient use, Decipher-MR supports a modular design that enables tuning of lightweight, task-specific decoders attached to a frozen pretrained encoder. Following this setting, we evaluate Decipher-MR across disease classification, demographic prediction, anatomical localization, and cross-modal retrieval, demonstrating consistent improvements over existing foundation models and task-specific approaches. These results position Decipher-MR as a versatile foundation for MRI-based AI in clinical and research settings.
@@ 2.1 Decipher-MR Pretraining
Decipher-MR was pretrained on a large and diverse MRI dataset comprising 22,594 studies and 203,233 series, with corresponding radiology reports available for 20,658 studies. The dataset spans a broad age range (0 to 90 years), covers a wide spectrum of body regions and MRI protocols, and features scans acquired from multiple vendors, including GE, Siemens, Philips, Toshiba, etc. Detailed quantitative distributions are summarized in Fig. 1a. This diversity ensures robust representation across various imaging systems and clinical environments.
@@ 2.2 Decipher-MR Learns Robust Visual Representation for Classification Tasks
A broad range of clinical and radiological tasks involving MRI can be formulated as classification problems, including disease diagnosis, demographic prediction, body region identification, image sequence classification, and contrast detection. We benchmarked our method against state-of-the-art foundation models using a simple probing setup, where the pretrained encoder was frozen and only a lightweight three-layer MLP attached on top was fine-tuned across these varied tasks. Our model consistently outperformed DINOv2[31] and BiomedCLIP[47] by a large margin, and demonstrated clear advantages over MedImageInsight[6], achieving average gains of 2.9% in disease classification (Supplementary Table 3), 3.0% in demographic prediction (Supplementary Table 4), and 0.2% in imaging-related attribute detection (Supplementary Table 5; Fig. 2a). The statistical significance of differences across tasks is reported in Supplementary Table 6. Performance differences between models became more pronounced under limited data settings (Supplementary Table 3-5), especially in tasks such as imaging attribute detection where performance otherwise tends to saturate (Fig. 2c). Further details on the classification tasks within each category, along with the datasets used, are provided in Methods Section 4.10 and Supplementary Table 2.
@@
Importantly, most of evaluations were performed without task-specific preprocessing (e.g., skull stripping, registration, or organ cropping) or fine-tuning, ensuring fair and direct model comparisons. Further performance improvements are possible through specialized model architectures and task-optimized data preprocessing. For instance, simply resizing MR images in-plane to match MedImageInsight’s input resolution substantially improved performance in tasks such as liver and prostate lesion classification, brain age regression, and attribute detection (Fig. 2a). Additional cropping around the region of interest can further enhance performance for localized tasks—for instance, prostate lesion classification improved by 7% when cropping around the prostate (ProstateCrop vs. Prostate in Fig.2a and Supplementary Table 3), further amplifying Decipher-MR’s advantage over other models. However, such cropping can impair performance on tasks that rely on broader anatomical context, such as age prediction from prostate MRI (Supplementary Table 3), which benefits from information beyond the prostate region.
@@ 2.3 Decipher-MR Benefits from Text and Diversity in Pretraining Data
Performance gains were observed with second-stage image–text contrastive pretraining, which outperformed models trained solely on imaging data (Fig. 2b) on most tasks, with significant gains in disease classification, brain age/sex prediction, and body region classification (Supplementary Table 6)). These results highlight the benefit of weak supervision from radiology reports and demonstrate the effectiveness of our two-stage pretraining strategy. Notable improvements were found in cardiac disease classification (+5.0%), prostate lesion classification (+2.4%), and several low-data scenarios, including brain age regression (+3.2%), body region detection (+4.2%), and contrast detection (+1.6%), with the latter two particularly pronounced under low-data scenarios (Fig. 2c).
@@ 2.4 Decipher-MR Supports Cross-modality Retrieval
On in-domain datasets, querying with either full reports or only the conclusion sections (Supplementary Fig. 1) retrieved the exact paired scan within the top 10 results in approximately 26% of cases, out of around 25,000 candidate MRI images from 2,500 studies, substantially outperforming MedImageInsight (5.1%) (Supplementary Table 8).
@@ 2.5 Decipher-MR Enables Robust Medical Image Segmentation with Fast Convergence
Across all datasets, Decipher-MR paired with a ResNet-based decoder (Method 4.12) consistently outperformed SAMMed3D[40] under the same frozen-encoder setting (Fig. 4a). Remarkably, even though only the decoder was trained, the performance was comparable to nnUNet[19], a widely used, fully tuned, state-of-the-art end-to-end segmentation framework in medical imaging (Fig. 4a). Qualitative examples in Fig. 4c show accurate segmentation of small abdominal organs and cardiac structures at both end-diastole (ED) and end-systole (ES), closely aligned with ground truth annotations. These results underscore the strong representational power of Decipher-MR across diverse anatomical regions.
@@ 2.6 Decipher-MR Supports Anomaly Localization and Visual Grounding
Beyond classification and organ segmentation, accurate pathology or anomaly localization is critical for the clinical deployment of AI models, especially when integrating human input. To investigate this, we evaluated Decipher-MR on anomaly localization tasks involving head and neck tumors and surgically removed organs (e.g., prostatectomy, hepatectomy, and pneumonectomy) (Fig. 4e), under two settings: standalone localization and visual grounding guided by human-provided text input describing anatomical location and pathology types. For each setting, we adopted state-of-the-art task-specific frameworks: DETR3D[43] for localization and MedRPG[5] for visual grounding, replacing and initializing their visual encoders with Decipher-MR (Method 4.13).
@@ 3 Discussion
In this study, we introduced Decipher-MR, an MRI-specific 3D vision-language foundation model pretrained on a large and diverse collection of MR datasets. We evaluated its capabilities across a comprehensive range of downstream tasks—including classification, cross-modal retrieval, organ segmentation, and anomaly localization—spanning multiple stages of the clinical workflow. Experimental results on both propriety and public datasets demonstrated that Decipher-MR is a powerful feature extractor for MR images, with competitive inference speed (Supplementary Table 11). It consistently outperformed other foundation models on simple probing tasks and achieved performance on par with or exceeding state-of-the-art, task-specific models trained end-to-end. One particularly interesting finding is that Decipher-MR maintains strong performance even though its pretraining data primarily consists of individuals under the age of 70 and rarely include cases of Alzheimer’s disease (six within 299 headneck or brain-scan patients aged 65 had reports referencing cognitive impairment or dementia.) [36]. Despite this, it performs competitively on related tasks, suggesting a degree of generalization beyond the original pretraining distribution.
@@
Despite these advantages, several limitations remain. First, we used Llama 3.3 to extract image-related information from radiology reports. Random human checks showed the task was straightforward and reliable, though full verification would further ensure consistency. Considering the human effort required, future validation could be performed using other LLMs as cross-checkers or judges. Second, although Decipher-MR showed strong performance in cross-modal retrieval, particularly for anatomical region retrieval, its effectiveness in pathology-focused retrieval was more limited. This likely stems from insufficient diversity in textual report data during pretraining: our reports came from a limited set of sources, where radiologists use highly standardized or repetitive terminology for similar conditions or negative findings, leading to many near-duplicated sentences and reduced generalization to more diverse texts. Increasing the diversity of pretraining data sources and incorporating LLM-based text augmentation to enrich report phrasing are potential directions to mitigate this. Third, we did not observe significant improvements in imaging attribute classification (e.g., sequence or contrast type) using additional text supervision. Incorporating DICOM metadata during pretraining may improve model capability in these areas and enable more radiology-relevant search functionalities. Moreover, larger gains from the second-stage text–image alignment may require incorporating localization-aware supervision that aligns textual descriptions with specific regions, rather than treating the entire image holistically. Such region-level alignment could improve the model’s ability to localize anatomical structures and pathological findings. Finally, while our frozen encoder setup has shown robust performance, encoder fine-tuning may be necessary to unlock further gains for certain applications. With the idea of maintaining a universal backbone, we plan to explore tuning strategies like LoRA and adapters to balance efficiency and task-specific performance.
@@ 4.2 Evaluation Dataset
Hold-out test set of our pretraining datasets: Contains around 2,000 studies and 20,000 MRI series, with radiology reports for most studies. MRI sequence details, body regions, and contrast usage were extracted from DICOM headers.
@@
ADNI: A dataset for Alzheimer’s disease research, containing diagnoses for cognitively normal (CN), mild cognitive impairment (MCI), and Alzheimer’s disease (AD) individuals, along with patient demographics. We used 1,732 baseline T1-weighted MRIs (339 AD, 871 MCI, and 522 CN) from ADNI1 and ADNI2 for classification evaluations.[33]
PICAI: A prostate cancer dataset comprising 1,500 bi-parametric MRI (bpMRI) exams per patient, including T2-weighted (T2W), high b-value diffusion-weighted imaging (DWI), and apparent diffusion coefficient (ADC) maps, along with annotated prostate and tumor masks and diagnostic labels (benign (1075 cases) vs. malignant (425 cases)) [35].
ACDC: A cardiac cine-MRI dataset with 150 exams across five subgroups, including four pathological conditions (myocardial infarction, dilated cardiomyopathy, hypertrophic cardiomyopathy, abnormal right ventricle) and one healthy group with 30 cases per category, along with segmentation masks for three heart regions [2].
LLD-MMRI: A multi-phase liver MRI dataset for lesion diagnosis, covering seven lesion types: three malignant (intrahepatic cholangiocarcinoma (58 cases), liver metastases (51 cases), hepatocellular carcinoma (157 cases)) and four benign (hepatic hemangioma (79 cases), hepatic abscess (54 cases), hepatic cysts (53 cases), focal nodular hyperplasia (46 cases)) [23].
MRART: A dataset comprising 438 T1-weighted structural MRI head images collected from 148 subjects, capturing three levels of motion artifacts: level 1 (129 images), level 2 (109 images), and level 3 (198 images) [29].
@@ 4.3 Model Architecture
For Decipher-MR, we employed a 3D Vision Transformer (ViT) [12] as the image encoder, using non-overlapping 8×8×8 volumetric patches. Each patch was flattened and passed through a linear projection to generate patch embeddings. Trilinear position embedding interpolation enables the model to handle inputs of variable sizes during inference and downstream applications. The rest of the architecture follows the standard ViT-Base configuration, consisting of 86 million parameters and a 768-dimensional embedding size; and can be found in the original ViT paper [12].
@@ 4.4 Model Pretraining
In the second stage of pretraining, we leveraged the pretrained vision and text encoders from the first stage and further refined them using an Image-Report Contrastive Learning approach [34] to align image and text embeddings in a shared multi-modal embedding space. During training, MRI-report pairs are processed to generate their respective embeddings. The model encourages embeddings from matching image-report pairs to be close in the shared space while pushing apart embeddings from non-matching pairs. Details regarding the construction of image and text inputs, along with the data sampling strategy designed to enhance contrastive learning, are provided in Section 4.6.
4.5 Data Augmentation
To enable a generalist foundation model capable of handling inconsistencies in MRI resolution, size, and acquisition protocols across scanners and body regions, we tailored our vision encoder training to be robust to such inconsistencies. Beyond standard augmentations like rotation and flipping, we customized the cropping strategy used in DINO’s teacher network. DINOv2 includes two global crops, and we adapted them for MRI by defining one crop that covers the entire scan resized to the input dimensions, and another partial crop extracted at a randomly selected resolution (0.5–1.5 mm in-plane, 1–6 mm out-of-plane) to introduce scale diversity, allowing the same scan to be seen at multiple resolutions. From each of these two global crops, we generated local crops with the same resolutions for the student network, enabling the model to learn consistent image-level representations. In the second stage of multi-modal contrastive training, we preserved more complete contextual information while keeping robustness to resolution variation by randomly cropping within 75%–100% of the image size and resizing to a relatively larger input size to retain fine details and resolutions (Section 4.7).
4.6 Organ-Based Study-level Data Sampling
@@ 4.7 Pretraining Data Processing
MRI Processing: we first applied foreground cropping to all MRI images using Otsu’s method with 256 bins to remove the black background. Each image was then rescaled to the range of -1 to 1 based on the 1st and 99th percentiles of pixel intensities. For the first-stage pretraining, we set the global crop size to 256×256×24 and the local crop size to 128×128×16, selected based on the median dimensions of our pretraining dataset. As described in Section 4.5, the first global crop was resized to match the target size while preserving the x-y aspect ratio, with black padding added as needed The second global crop was a random crop with variable resolution. In the second-stage contrastive training, to capture a larger portion of the input image while minimizing information loss from resizing, we increased the crop size to 384×384×48.
Report Processing: we first used Llama 3.3 [14] to process the noisy MRI reports by removing study details unrelated to imaging and extracting organ-specific descriptions and conclusions into separate sections, including any identified abnormalities. We then mapped each organ section to a list of standard SNOMED anatomical entities [22] using AWS Comprehend Medical [1] for the sampling process described in Section 4.6. We randomly inspected on average 30 processed reports per body region and refine the prompts to enhance robustness of the processing pipeline. Additional details, including the prompt used and examples of reports before and after processing, are provided in Supplementary Fig. 1.